Allison and joe have yo share $1715 in the ratio 3:4

How much does joe get


1
Expert's answer
2021-10-24T17:28:13-0400

If they share $1715 in the ratio 3:4, then 1 part is 1715/7 = $245.

Allison will receive 245×3 = $735.

Joe will receive 245×4 = $980.